Teal Inc. owns and operates a number of hardware stores in the New England region. Recently, the company has decided to locate another store in a rapidly growing area of Maryland. The company is trying to decide whether to purchase or lease the building and related facilities.

Purchase: The company can purchase the site, construct the building, and purchase all store fixtures. The cost would be $1,867,900. An immediate down payment of $402,100 is required, and the remaining $1,465,800 would be paid off over 5 years at $351,600 per year (including interest payments made at end of year). The property is expected to have a useful life of 12 years, and then it will be sold for $510,600. As the owner of the property, the company will have the following out-of-pocket expenses each period.

Property taxes (to be paid at the end of each year)

$40,790

Insurance (to be paid at the beginning of each year)

27,310

Other (primarily maintenance which occurs at the end of each year)

16,330

$84,430

Lease: First National Bank has agreed to purchase the site, construct the building, and install the appropriate fixtures for Teal Inc. if Teal will lease the completed facility for 12 years. The annual costs for the lease would be $263,970. Teal would have no responsibility related to the facility over the 12 years. The terms of the lease are that Teal would be required to make 12 annual payments (the first payment to be made at the time the store opens and then each following year). In addition, a deposit of $107,900 is required when the store is opened. This deposit will be returned at the end of the 12th year, assuming no unusual damage to the building structure or fixtures.

Compute the present value of lease vs purchase. (Currently, the cost of funds for Teal Inc. is 11%.) (Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 458,581.)
Answer #1

Answer:

Purchase
a b c d e=a+b+c+d f g=e*f
Year Down Payment, Installment & Sale   Property tax year end Insurance at year beginning Other maintainence year end Total cash outflow Present value factor@11% Present value of cashoutflow
0               4,02,100              27,310            4,29,410 1 1 429410
1               3,51,600                    40,790              27,310              16,330            4,36,030 0.90090 =(1/1.11)^1                  3,92,820
2               3,51,600                    40,790              27,310              16,330            4,36,030 0.81162 =(1/1.11)^2                  3,53,892
3               3,51,600                    40,790              27,310              16,330            4,36,030 0.73119 =(1/1.1)^3                  3,18,821
4               3,51,600                    40,790              27,310              16,330            4,36,030 0.65873 =(1/1.11)^4                  2,87,226
5               3,51,600                    40,790              27,310              16,330            4,36,030 0.59345 =(1/1.11)^5                  2,58,763
6                    40,790              27,310              16,330               84,430 0.53464 =(1/1.11)^6                     45,140
7                    40,790              27,310              16,330               84,430 0.48166 =(1/1.11)^7                     40,666
8                    40,790              27,310              16,330               84,430 0.43393 =(1/1.11)^8                     36,636
9                    40,790              27,310              16,330               84,430 0.39092 =(1/1.11)^9                     33,006
10                    40,790              27,310              16,330               84,430 0.35218 =(1/1.11)^10                     29,735
11                    40,790              27,310              16,330               84,430 0.31728 =(1/1.11)^11                     26,788
12              -5,10,600                    40,790              27,310              16,330          -4,26,170 0.28584 =(1/1.11)^12                 -1,21,817
Present cost of purchasing                21,31,087
Leasing
a b c=a+b d e=c*d
Year Annual lease at year beginning Deposit Total cash outflow Present value factor@11% Present value of cashoutflow
0               2,63,970                 1,07,900           3,71,870 1 1 371870
1               2,63,970           2,63,970 0.90090 =(1/1.11)^1          2,37,811
2               2,63,970           2,63,970 0.81162 =(1/1.11)^2          2,14,244
3               2,63,970           2,63,970 0.73119 =(1/1.1)^3          1,93,013
4               2,63,970           2,63,970 0.65873 =(1/1.11)^4          1,73,885
5               2,63,970           2,63,970 0.59345 =(1/1.11)^5          1,56,653
6               2,63,970           2,63,970 0.53464 =(1/1.11)^6          1,41,129
7               2,63,970           2,63,970 0.48166 =(1/1.11)^7          1,27,143
8               2,63,970           2,63,970 0.43393 =(1/1.11)^8          1,14,544
9               2,63,970           2,63,970 0.39092 =(1/1.11)^9          1,03,192
10               2,63,970           2,63,970 0.35218 =(1/1.11)^10             92,966
11               2,63,970           2,63,970 0.31728 =(1/1.11)^11             83,753
12                -1,07,900          -1,07,900 0.28584 =(1/1.11)^12            -30,842
Present cost of leasing        19,79,362
Since the cost of leasing is less than cost of purchasing. Lease should be opted.